# Modeling and Beamforming Optimization for Pinching-Antenna Systems

### v2

@@ Abstract
Figure 1: Comparison between conventional wireless system (left) and PASS (right).
@@ I Introduction
Pinching-Antenna SyStem (PASS) is a revolutionary technique for addressing the challenges of free-space pathloss and LoS blockage encountered by conventional multi-antenna technologies. This technique was originally proposed and prototyped by NTT DOCOMO in 2022 [15]. As illustrated on the right of Fig. 1, PASS employs a dielectric waveguide as its primary transmission medium, which is known for its exceptionally low propagation loss (e.g., 0.01 dB/m [16]). By pinching a small separated dielectric element, referred to as a pinching antenna, onto the waveguide, the system enables signal emission from the waveguide into the pinching antenna, which then radiates the signal into free space. Building on this principle, waveguides can be pre-deployed to extend service coverage, allowing pinching antennas to be placed at positions close to users. This strategic placement transforms the wireless system into a near-wired system and hence establishes strong LoS links with users, effectively minimizing free-space path loss and mitigating blockage issues. Additionally, unlike existing MIMO systems, PASS allows both the number and positions of pinching antennas to be easily adjusted by simply pinching them to or releasing them from the waveguide [15]. This feature provides a low-cost and scalable approach to implementing MIMO while also facilitating the so-called pinching beamforming, which enhances communication performance by dynamically optimizing antenna positions [17].
Given the successful prototyping of PASS by NTT DOCOMO, theoretical research on this topic has been steadily growing, though it remains in its early stages. The first theoretical study on PASS for the communication system design was presented in [18], where the authors provided a comprehensive analysis and developed low-complexity pinching beamforming designs for fundamental single-user and two-user scenarios. The array gain achieved by multiple pinching antennas on a waveguide was analyzed in [19], unveiling the optimal number of antennas and their spacing for maximizing the beamforming gain. The authors of [20] explored a downlink PASS architecture utilizing multiple waveguides, each equipped with a single pinching antenna, and proposed a greedy approach for jointly optimizing the transmit and pinching beamforming. Meanwhile, [21] examined a more generalized scenario, where multiple pinching antennas were deployed on each waveguide, and introduced a graph neural network (GNN)-based deep learning method to address the corresponding joint beamforming optimization problem.
@@
  • We derive a novel signal model for PASS based on the proposed physics framework, revealing the inherent coupling effect between the radiation power of pinching antennas deployed on the same waveguide. Leveraging this coupling relationship, we introduce two simplified power models and their respective implementation methods: equal power and proportional power models.
@@ II-A Physics-based Hardware Model
The core principle of pinching antennas relies on the phenomenon where a portion of the EM waves propagating through a dielectric waveguide is induced into an adjacent dielectric material (i.e., a pinching antenna) when the two are placed in close proximity. To accurately model the signal radiated by the pinching antenna, it is essential to characterize the EM fields within both the waveguide and the pinching antenna, as well as to understand the interaction between these fields.
@@
From these expressions, it is evident that a maximum fraction of $\left(\kappa/\phi\right)^{2}$ of the total power can be transferred to the pinching antenna. In the special case where the waveguide and the pinching antenna have the same effective refractive index (i.e., $\beta_{\mathrm{g}}=\beta_{\mathrm{p}}$), we have $\Delta\beta=0$ and $\phi=\kappa$. Under this condition, the expressions for $A(x)$ and $B(x)$ can be simplified to:
@@
(Power Relationship) Based on the above physics model, the power of the signals within the waveguide and radiated by the pinching antenna is determined by the coupling length $L$. In the case of $\beta_{\mathrm{g}}=\beta_{\mathrm{p}}$, a complete power transfer to the pinching antenna becomes achievable, enabling full signal radiation, which is achieved when $L=\pi/(2\kappa)$ according to (12). However, this strategy is applicable for the case with a single pinching antennas, but not for the case with multiple pinching antennas on the same waveguide. For the latter case, the radiation power from each pinching antenna can be adjusted by modifying the coupling length $L$. In other words, the transmit powers of the multiple pinching antennas can be controlled, which leads to extra degrees of freedom for the system design; howerver, this would likely lead to active pinching antenna design and additional system complexity. Therefore, this paper focuses on a purely passive pinching antenna design with a preconfigured coupling length.
@@ II-B Signal Model
In (II-A) and (II-A), we derived the physical models describing the EM wave behavior in the PASS system. Building upon these foundational physics models, we can now formulate simplified signal models commonly employed in the wireless communication system design. In the following, we begin by considering the scenario with a single pinching antenna coupled to the waveguide and subsequently extend the model to cases involving multiple pinching antennas.
@@ II-B1 A Single Pinching Antenna
The radiated signal propagates through free space to reach the user, where attenuation due to free-space path loss must be considered, leading to the following signal model:
@@ II-B2 Multiple Pinching Antennas
(Pinching Beamforming) It can be observed from (23)-(25) that adjusting the positions of the pinching antennas, determined by $\mathbf{x}$, alters the phase and the large-scale path loss (amplitude) of the signal received by the user. Analogous to conventional multi-antenna beamforming, the signals from multiple pinching antennas can be combined either constructively or destructively at the user by carefully optimizing the positions of the pinching antennas. This new capability for signal reconfiguration introduced by PASS is referred to as pinching beamforming.
@@ III-C2 Subproblem With Respect to $\mathbf{U}$
This subproblem is non-convex due to the quadratic term $\left|\mathbf{u}_{k}^{H}\mathbf{v}_{k}\right|^{2}$ on the left-hand side of the inequality constraints. We exploit the successive convex approximation method to address it. Let $\mathbf{U}^{(t)}=\big{[}\mathbf{u}_{1}^{(t)},\dots,\mathbf{u}_{K}^{(t)}\big{]}$ denote the solution of $\mathbf{U}$ obtained in the previous iteration of the inner loop. Then, according to the first-order Taylor expansion, a low bound on the quadratic term $\left|\mathbf{u}_{k}^{H}\mathbf{v}_{k}\right|^{2}$ can be obtained as follows:
$\left|\mathbf{u}_{k}^{H}\mathbf{v}_{k}\right|^{2}$ $\geq-\left(\mathbf{u}_{k}^{(t)}\right)^{H}\mathbf{v}_{k}\mathbf{v}_{k}^{H}\mathbf{u}_{k}^{(t)}$
$+2\Re\left\{\left(\mathbf{u}_{k}^{(t)}\right)^{H}\mathbf{v}_{k}\mathbf{v}_{k}^{H}\mathbf{u}_{k}\right\}\triangleq h\left(\mathbf{u}_{k}|\mathbf{u}_{k}^{(t)}\right),$ (47)
where the equality is achieved at $\mathbf{u}_{k}=\mathbf{u}_{k}^{(t)}$. Therefore, problem (46) can be approximated by
@@
$\mathrm{s.t.}\quad$ $\frac{1}{\gamma_{k}}h\left(\mathbf{u}_{k}|\mathbf{u}_{k}^{(t)}\right)\geq\sum_{i=1,i\neq k}^{K}\left|\mathbf{u}_{k}^{H}\mathbf{v}_{i}\right|^{2}+\frac{\sigma_{k}^{2}}{P},\forall k.$ (48b)
@@ III-C4 Subproblem With Respect to $\mathbf{X}$
It can be observed that problem (55) is to find the minimum of a single-variable function over a finite interval. Consequently, it can be efficiently solved using a low-complexity one-dimensional search method, applicable to both continuous and discrete activation scenarios. For continuous activation, the feasible set $\mathcal{S}=[0,x_{\max}]$ is discretized into a fine grid. The optimal solution that minimizes $g(x_{n,m})$ is then searched within the sampled feasible set, subject to the constraint in (55b). For discrete activation, the optimal solution can be directly identified by searching over the discrete set of feasible values. Based on the above approaches, the overall element-wise algorithm for solving problem (52) is summarized in Algorithm 1. It is worth noting that the one-dimensional search effectively addresses the multimodality issue by preventing the optimization from getting stuck in local optima.
@@ IV Numerical Results
In this section, numerical results are presented to illustrate the advantages of PASS and validate the effectiveness of the proposed algorithms. Unless stated otherwise, the following setup is used throughout the simulations. We consider $N=5$ waveguides, each equipped with $M=6$ pinching antennas and fed by a dedicated RF chain, serving $K=4$ users. The geometric configuration of the setup is depicted in Fig. 5, where the key parameters are set as $d_{0}=15$ m, $d_{x}=30$ m, $d_{y}=3$ m, and $d_{z}=10$ m. The waveguides are uniformly deployed in parallel along the $x$-axis with a spacing of $6$ m, while the users are randomly positioned within the designated service area. The maximum range and minimum spacing of pinching antenna positions are set to $x_{\max}=50$ m and $\Delta x=0.1$ m, respectively. For discrete activation, the number of discrete positions is set to $10$ per meter.
@@
Furthermore, the carrier frequency, channel gain, noise power, and effective index of the waveguide are set to $15$ GHz, $|\eta|^{2}=\left(\frac{\lambda}{4\pi}\right)^{2}=-56$ dB, $\sigma_{k}^{2}=-80$ dBm, and $n_{g}=1.4$, respectively. The total power radiated from all pinching antennas along each waveguide is constrained by $\sum_{m=1}^{M}\alpha_{m}^{2}=0.9$, which applies to both equal and proportional power models. The minimum SINR of all users is set to $20$ dB. For the proposed algorithms, the penalty factor is initialized as $\rho=10$, the reduction factor is set to $\epsilon=0.1$, and the convergence threshold is set to $10^{-3}$. The number of search points in the one dimensional search for continuous activation is set to $10^{6}$. Furthermore, due to its sensitivity to initialization, the penalty-based method utilizes the antenna positions obtained from the ZF-based algorithm as its starting point. Unless otherwise specified, all subsequent results are obtained by averaging over $100$ random samples of user positions.
@@
(b) Constraint violation in Algorithm 2.
@@
  • Conventional MIMO: In this benchmark, a conventional MIMO BS is positioned at $(0,0,3)$ and equipped with a uniform linear array along $x$-axis comprising $N=5$ antennas, each connected to a dedicated RF chain. The antenna spacing is set to half the wavelength. Under this setup, the signal received by user $k$ is given by
@@
  • Massive MIMO: In this benchmark, we consider a massive MIMO BS positioned at $(0,0,3)$, equipped with a uniform linear array along $x$-axis. The number of antennas and RF chains is set to $N=30$ and $N_{\mathrm{RF}}=5$, respectively, matching the configuration of the considered PASS system. Furthermore, since each RF chain in the PASS system is connected to only a subset of overall pinching antennas, we assume a similar sub-connected hybrid beamforming architecture at the massive MIMO BS, where each RF chain is connected to a subset of antennas via phase shifters [28]. Under this setup, the signal received by user $k$ is given by
@@ IV-A Performance of the Proposed Algorithms
Fig. 6 illustrates the convergence behavior the proposed algorithms. Specifically, the convergence behavior of the transmit power $P$ in Algorithm 2 is shown in Fig. 6(a). It is interesting to observe that the transmit power $P$ is not reduced monotonically as the iteration progresses, but exhibits an oscillating upward trend before converging to a stable value. This phenomenon is due to the utilization of the penalty method. Specifically, in the initial iterations, when the penalty factor $\rho$ is large, the transmit power $P$ is minimized with an almost unconstrained auxiliary channel matrix $\mathbf{U}$, allowing it to converge to a very low stable value in the inner loop. However, as the penalty factor decreases in the outer loop, constraint violations gradually diminish as shown in Fig. 6(b), forcing the auxiliary channel matrix $\mathbf{U}$ to conform to the channel structure and constraints in PASS. Consequently, the transmit power increases as the outer loop progresses. Fig. 6(c) demonstrates the fast convergence of the proposed Algorithm 3. Despite its low complexity, Algorithm 3 achieves performance comparable to Algorithm 2 across various system setups, as shown in Fig. 7. Therefore, in the following simulations, only the results obtained by Algorithm 2 are presented to focus on the performance gains achieved by PASS.
Figure 8: Transmit power versus the minimum SINR.
Furthermore, Fig. 6 reveals that different values of the reduction factor $\epsilon$ affect the convergence speed. Specifically, the algorithm converges more quickly with a smaller $\epsilon$, while a larger $\epsilon$ leads to improved performance. However, it is worth noting that the performance gain from increasing $\epsilon$ from $0.1$ to $0.5$ is negligible. Therefore, we set $\epsilon=0.1$ for the remaining simulations.
IV-B Transmit Power Versus the Minimum SINR

### v3

@@ Abstract
Figure 1: Schematic illustration of the conventional wireless system (left) and the PASS (right).
@@ I Introduction
Pinching-Antenna SyStem (PASS) is a revolutionary technique for addressing the challenges of free-space pathloss and LoS blockage encountered by conventional multi-antenna technologies. This technique was originally proposed and prototyped by NTT DOCOMO in 2022 [15]. As illustrated on the right of Fig. 1, PASS employs a dielectric waveguide as its primary transmission medium, which is known for its exceptionally low propagation loss (e.g., 0.01 dB/m [16]). By pinching a small separated dielectric element, referred to as a pinching antenna, onto the waveguide, the system enables signal emission from the waveguide into the pinching antenna, which then radiates the signal into free space. Building on this principle, waveguides can be pre-deployed to extend service coverage, allowing pinching antennas to be placed at positions close to users. This strategic placement transforms the wireless system into a near-wired system and hence establishes strong LoS links with users, effectively minimizing free-space path loss and mitigating blockage issues, sharing a similar concept with surface wave communications where waves propagate over a slit along a metasurface and radiate from pre-defined locations [17]. These unique advantages make PASS a promising alternative for indoor communications. Additionally, unlike existing MIMO systems and surface wave communications, PASS allows both the number and positions of pinching antennas (i.e., the radiation point) to be easily adjusted by simply pinching them to or releasing them from the waveguide [15]. This feature provides a low-cost and scalable approach to implementing MIMO while also facilitating the so-called pinching beamforming, which enhances communication performance by dynamically optimizing antenna positions [18].
Given the successful prototyping of PASS by NTT DOCOMO, theoretical research on this topic has been steadily growing, though it remains in its early stages. The first theoretical study on PASS for the communication system design was presented in [19], where the authors provided a comprehensive analysis and developed low-complexity pinching beamforming designs for fundamental single-user and two-user scenarios. The array gain achieved by multiple pinching antennas on a waveguide was analyzed in [20], unveiling the optimal number of antennas and their spacing for maximizing the beamforming gain. The authors of [21] explored a downlink PASS architecture utilizing multiple waveguides, each equipped with a single pinching antenna, and proposed a greedy approach for jointly optimizing the transmit and pinching beamforming. Meanwhile, [22] examined a more generalized scenario, where multiple pinching antennas were deployed on each waveguide, and introduced a graph neural network-based deep learning method to address the corresponding joint beamforming optimization problem.
@@
  • We derive a novel signal model for PASS based on the proposed physics framework, assuming identical effective refractive indices for the waveguides and pinching antennas. This model reveals the inherent coupling effect between the radiation power of pinching antennas deployed on the same waveguide. Leveraging this coupling relationship, we introduce two simplified power models and their respective implementation methods: equal power and proportional power models.
@@ II-A Physics-based Hardware Model
The core principle of pinching antennas relies on the phenomenon where a portion of the EM waves propagating through a dielectric waveguide is induced into an adjacent dielectric material (i.e., a pinching antenna) when the two are placed in close proximity [^1] It is possible to apply pinching antennas on both sides of a specific location along the waveguide, which, however, would significantly increase the complexity of the model. In this paper, we restrict our analysis to a single pinching antenna per location to maintain a clear and tractable framework.. To accurately model the signal radiated by the pinching antenna, it is essential to characterize the EM fields within both the waveguide and the pinching antenna, as well as to understand the interaction between these fields.
@@
From these expressions, it is evident that a maximum fraction of $\left(\kappa/\phi\right)^{2}$ of the total power can be transferred to the pinching antenna. In the rest of this paper, we consider a special case where the waveguide and the pinching antenna have the same effective refractive index (i.e., $\beta_{\mathrm{g}}=\beta_{\mathrm{p}}$), where we have $\Delta\beta=0$ and $\phi=\kappa$. Under these conditions, the expressions for $A(x)$ and $B(x)$ can be simplified to:
@@
(Power Relationship) Based on the above physics model, the power of the signals within the waveguide and radiated by the pinching antenna is determined by the coupling length $L$. In the case of $\beta_{\mathrm{g}}=\beta_{\mathrm{p}}$, a complete power transfer to the pinching antenna becomes achievable, enabling full signal radiation, which is achieved when $L=\pi/(2\kappa)$ according to (12). However, this strategy is applicable for the case with a single pinching antennas, but not for the case with multiple pinching antennas on the same waveguide. For the latter case, the radiation power from each pinching antenna can be adjusted by modifying the coupling length $L$. In other words, the transmit powers of the multiple pinching antennas can be controlled, which leads to extra degrees of freedom (DoFs) for the system design; howerver, this would likely lead to active pinching antenna design and additional system complexity. Therefore, this paper focuses on a purely passive pinching antenna design with a preconfigured coupling length.
@@ II-B Signal Model
In (II-A) and (II-A), we derived the physical models describing the EM wave behavior in the PASS system, assuming identical effective refractive indices for the waveguides and pinching antennas. Building upon these foundational physics models, we can now formulate simplified signal models commonly employed in the wireless communication system design. In the following, we begin by considering the scenario with a single pinching antenna coupled to the waveguide and subsequently extend the model to cases involving multiple pinching antennas.
@@ II-B1 A Single Pinching Antenna
The radiated signal propagates through free space to reach the user, where attenuation due to free-space path loss must be considered, leading to the following signal model [^2] In our signal model, we consider only LoS path in the free space due to two reasons. First, it is important to note that dielectric waveguides typically operate in high-frequency bands, such as the millimeter-wave bands [15]. At these high frequencies, the LoS path can be 20 dB stronger than NLoS paths [25], making the impact of NLoS components almost negligible. Second, even if NLoS paths are relevant, their path loss exponents are generally higher than that of the LoS path. Thus, the performance benefits provided by PASS can be even more significant, since it effectively reduce the transmission distance.:
@@ II-B2 Multiple Pinching Antennas
(Pinching Beamforming) Pinching antennas extract signals from waveguides through EM coupling with no physical connection to the waveguide. Therefore, it is possible to dynamically reposition the pinching antennas. For instance, a pinching antenna can be made movable by placing it on a pre-installed track parallel to the waveguide, which is similar to how a motorized camera slider operates [19]. It can be observed from (23)-(25) that adjusting the positions of the pinching antennas, determined by $\mathbf{x}$, alters the phase and the large-scale path loss (amplitude) of the signal received by the user. Analogous to conventional multi-antenna beamforming, the signals from multiple pinching antennas can be combined either constructively or destructively at the user by carefully optimizing the positions of the pinching antennas. This new capability for signal reconfiguration introduced by PASS is referred to as pinching beamforming.
@@ III-C2 Subproblem With Respect to $\mathbf{U}$
This subproblem is non-convex due to the quadratic term $\left|\mathbf{u}_{k}^{H}\mathbf{v}_{k}\right|^{2}$ on the left-hand side of the inequality constraints. We exploit the successive convex approximation method to address it. Let $\mathbf{U}^{t}=\big{[}\mathbf{u}_{1}^{t},\dots,\mathbf{u}_{k}^{t}\big{]}$ denote the solution of $\mathbf{U}$ obtained in the previous iteration of the inner loop. Then, according to the first-order Taylor expansion, a low bound on the quadratic term $\left|\mathbf{u}_{k}^{H}\mathbf{v}_{k}\right|^{2}$ can be obtained as follows:
$\left|\mathbf{u}_{k}^{H}\mathbf{v}_{k}\right|^{2}$ $\geq-\left(\mathbf{u}_{k}^{t}\right)^{H}\mathbf{v}_{k}\mathbf{v}_{k}^{H}\mathbf{u}_{k}^{t}$
$+2\Re\left\{\left(\mathbf{u}_{k}^{t}\right)^{H}\mathbf{v}_{k}\mathbf{v}_{k}^{H}\mathbf{u}_{k}\right\}\triangleq h\left(\mathbf{u}_{k}|\mathbf{u}_{k}^{t}\right),$ (47)
where the equality is achieved at $\mathbf{u}_{k}=\mathbf{u}_{k}^{t}$. Therefore, problem (46) can be approximated by
@@
$\mathrm{s.t.}\quad$ $\frac{1}{\gamma_{k}}h\left(\mathbf{u}_{k}|\mathbf{u}_{k}^{t}\right)\geq\sum_{i=1,i\neq k}^{K}\left|\mathbf{u}_{k}^{H}\mathbf{v}_{i}\right|^{2}+\frac{\sigma_{k}^{2}}{P},\forall k.$ (48b)
@@ III-C4 Subproblem With Respect to $\mathbf{X}$
It can be observed that problem (55) is to find the minimum of a single-variable function over a finite interval. Depending on whether the activation is continuous or discrete, the solution can be obtained as follows:
  • For continuous activation, to solve the multimodal objective function, an efficient approach is one-dimensional grid search [29], which prevents the optimization from getting stuck in local optima. In particular, the feasible set $\mathcal{S}=[0,x_{\max}]$ is discretized into a fine grid. The optimal solution that minimizes $g(x_{n,m})$ is then searched over the fine grid subject to the constraint (55b). To mitigate off-grid errors, a gradient-based refinement (e.g., Newton’s method) can be applied starting from the grid-search solution, which, however, additional higher computational complexity.
  • For discrete activation, the optimal solution can be directly identified by searching over the discrete feasible set $\mathcal{S}=\{0,\frac{x_{\max}}{Q-1},\dots,x_{\max}\}$, subject to the constraint (55b).
Based on the above discussion, both continuous and discrete positions of pinching antenna can be optimized through a one-dimensional search. Therefore, the overall element-wise algorithm for solving problem (52) is summarized in Algorithm 1.
@@ III-C5 Overall Algorithm
Although Algorithm 2 exhibits relatively high computational complexity, it possesses two important advantages. On the one hand, it directly addresses the original problem formulation, and the penalty-based method employed guarantees convergence to a stationary point [27, 28]. Consequently, Algorithm 2 serves as a reliable baseline for evaluating other approaches, such as the zero-forcing algorithm presented in the subsequent section. On the other hand, Algorithm 2 is highly generalizable, as it addresses coupling issues caused by pinching antennas by optimizing penalty terms independently of specific objective functions. Thus, it can be extended to address other optimization problems arising in pinching-antenna designs.
@@ IV Numerical Results
In this section, numerical results are presented to illustrate the advantages of PASS and validate the effectiveness of the proposed algorithms. Unless stated otherwise, we consider an indoor communication scenario, and the following setup is used throughout the simulations. We consider $N=5$ waveguides, each equipped with $M=6$ pinching antennas and fed by a dedicated RF chain, serving $K=4$ users. The geometric configuration of the setup is depicted in Fig. 5, where the key parameters are set as $d_{0}=15$ m, $d_{x}=30$ m, $d_{y}=3$ m, and $d_{z}=10$ m. The waveguides are uniformly deployed in parallel along the $x$-axis with a spacing of $6$ m, while the users are randomly positioned within the designated service area. The maximum range and minimum spacing of pinching antenna positions are set to $x_{\max}=50$ m and $\Delta x=0.1$ m, respectively. For discrete activation, the number of discrete positions is set to $10$ per meter.
@@
Furthermore, the carrier frequency, channel gain, noise power, and effective index of the waveguide are set to $15$ GHz, $|\eta|^{2}=\left(\frac{\lambda}{4\pi}\right)^{2}=-56$ dB, $\sigma_{k}^{2}=-80$ dBm, and $n_{g}=1.4$, respectively. The total power radiated from all pinching antennas along each waveguide is constrained by $\sum_{m=1}^{M}\alpha_{m}^{2}=0.9$, which applies to both equal and proportional power models. The minimum SINR of all users is set to $\gamma_{k}=20$ dB. For the proposed algorithms, the penalty factor is initialized as $\rho=10$, the reduction factor is set to $\epsilon=0.1$, and the convergence threshold is set to $10^{-3}$. The number of search points in the one dimensional search for continuous activation is set to $10^{6}$. Furthermore, due to its sensitivity to initialization, the penalty-based method utilizes the antenna positions obtained from the ZF-based algorithm as its starting point. Unless otherwise specified, all subsequent results are obtained by averaging over $100$ random samples of user positions.
@@
(b) Constraint violation (57) in Algorithm 2.
@@
  • Conventional MIMO: In this benchmark, a conventional MIMO BS is positioned at $(0,0,3)$ and equipped with a uniform linear array along $x$-axis comprising $N_{\mathrm{RF}}=5$ chains, each connected to a single antenna. The antenna spacing is set to half the wavelength [^3] Movable antennas have recently garnered significant attention. Their primary advantage lies in enhancing spatial DoFs when only a limited number of antennas are deployed within a given aperture. However, in our baselines, we consider a dense deployment with half-wavelength spacing, which is known to fully exploit the available spatial DoFs [31]. Therefore, we assume the antennas to be fixed and non-movable in our baselines.. Under this setup, the signal received by user $k$ is given by
@@
  • Massive MIMO: In this benchmark, we consider a massive MIMO BS positioned at $(0,0,3)$, equipped with a uniform linear array with half-wavelength spacing along $x$-axis. The number of antennas and RF chains is set to $N=30$ and $N_{\mathrm{RF}}=5$, respectively, matching the configuration of the considered PASS system. Furthermore, since each RF chain in the PASS system is connected to only a subset of overall pinching antennas, we assume a similar sub-connected hybrid beamforming architecture at the massive MIMO BS, where each RF chain is connected to a subset of antennas via phase shifters [32]. Under this setup, the signal received by user $k$ is given by
@@ IV-A Performance of the Proposed Algorithms
Fig. 6 illustrates the convergence behavior the proposed algorithms. Specifically, the convergence behavior of the transmit power $P$ in Algorithm 2 is shown in Fig. 6(a). It is interesting to observe that the transmit power $P$ is not reduced monotonically as the iteration progresses, but exhibits an oscillating upward trend before converging to a stable value. This phenomenon is due to the utilization of the penalty method. Specifically, in the initial iterations, when the penalty factor $\rho$ is large, the transmit power $P$ is minimized with an almost unconstrained auxiliary channel matrix $\mathbf{U}$, allowing it to converge to a very low stable value in the inner loop. However, as the penalty factor decreases in the outer loop, constraint violations gradually diminish as shown in Fig. 6(b), forcing the auxiliary channel matrix $\mathbf{U}$ to conform to the channel structure and constraints in PASS. Consequently, the transmit power increases as the outer loop progresses. Fig. 6(c) demonstrates the fast convergence of the proposed Algorithm 3. Despite its low complexity, Algorithm 3 achieves performance comparable to Algorithm 2 across various system setups, as shown in Fig. 7.
Furthermore, Figs. 6(a) and 6(b) also reveal that different values of the reduction factor $\epsilon$ affect the convergence speed. Specifically, the algorithm converges more quickly with a smaller $\epsilon$, while a larger $\epsilon$ leads to improved performance by allowing the algorithm to more thoroughly minimize transmit power before reducing the penalty terms. However, it is worth noting that the performance gain from increasing $\epsilon$ from $0.1$ to $0.5$ is negligible. In the remaining simulations, only the results obtained by Algorithm 2 are presented to focus on the performance gains achieved by PASS.
IV-B Impact of the Minimum SINR